calculate the number of moles in 1.00 * 10(-3) g of CuCl2
1
Expert's answer
2013-02-20T09:52:10-0500
For answering this question you need to use one simply formula,that is used for finding amount (or number of moles)

The formula is n=m/Mw,

where n is amount
m is mass
& Mw is molecular mass

For CuCl2& m = 1.00 * 10(-3) and Mw = 134.452

n = 1.00 * 10(-3) / 134.452 = 7.44*10(-6) mol